A cob Douglas production function for a firm is given as Q=4L ½K½. The firm has also established that wage rate and interest paid on capital are $3 and $5 respectively for a production period. The firm intents to spend $200 million for the period on production cost. Compute the levels of capital and labor that will maximize output. What is the maximum output?    


1
Expert's answer
2021-03-10T07:23:36-0500

Form a Lagragian equation

Q=4L0.5K0.5 Subject to wL+rK=C

L=4L0.5K0.5 -"\lambda"(wL+rK-C)

"\delta"L/"\delta"L=2L-0.5K0.5-"\lambda"w=0..........(i)

"\delta"L/"\delta"K=2L0.5K-0.5-"\lambda"r=0...........(ii)

"\delta"L/"\delta""\lambda"=wL+rK-C=0.................(iii)

Divide equation (i) and (ii)

K/L=w/r and thus K=wL/r and L=rK/w

Replacing the two equation on equation (iii)

w(kr/w)+rK=C thus K*=C/2r

K*=200/2*5 =20

wL+r(wL/r)=C

wL+wL=C

L*=C/2w

L*=200/2*3 = 33.33

Hence the optimum output is;

Q=4(33.33)0.5(20)0.5

Q=4(4.58)(4.5)

Q=104.40 units - Maximum Output.